The software group continues to test different aspects of the TCS. They found many of the bugs that stopped us from generating trajectories last week. We should now be able to generate trajectories at any Az and for normal sidereal objects and for geostationary satellites (useful for diagnostics). We do
continue to run into some small teathing problems with the new tracker as we start running it through its paces, e.g. loose connectors, cables that are not routed the correct way and blown fuses.
The electrical and mechanical teams worked in the Krabbenhut this week and was able to remove the old Champion compressor, install new vertical air tank and move the Power-X compressor which supplies the instrument air. This was all done to make room for the Nitrogen Vaporizor which will generate very pure N2 gas for instrument purging from our large LN tank.
Our electronics tech continues to work with the Ultra-pure water system trying to add reliability and consistency to the system. He can now produce 17.1 Mohms water (any reading above 10 Mohms is considered is considered Ultra High Purity for water, with 18.2 Mohms being the ultimate maximum).
The mirror team finished the piston work this week and we hope to have the next round of piston be done with the DMI rather than with the sphereometer.
In the Facility Thermal Management Project Veliz Construction completed the welding of pylon frames (to support ducting) together in place over side of the mountain and ARC is working on the ducting next to the Krabbenhut and down the side of the mountain. ARC mechanical has completed the WFU glycol manifold at the outside of the dome. This is the point where HET personnel will continue the piping layout into the dome.